OpenStudy (anonymous):

Write an equation for the translation of y=6/x that has the asymptotes x=4 and y=5

OpenStudy (anonymous):

@GeraldJRJunior @mathmale @HourglassMage @MathsPro

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you please help me @Hotchellerae21 ?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (anonymous):

So where do we start?

OpenStudy (anonymous):

y=6/x

OpenStudy (anonymous):

Okay, then what?

OpenStudy (anonymous):

i have a link that can help us can i use it

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

let me pull it up kk

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

@Hotchellerae21 you still there?

OpenStudy (anonymous):

ya waiting for it to load

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

it is not loading

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

with out that i can't help

OpenStudy (anonymous):

Ok, but what is your best guess?

OpenStudy (e.mccormick):

The fact it asks for a translation means it will have some form of: \(y=\dfrac{6}{x-h}+k\) where h is the shift in x and k is the shift in y. However, it identifies the shifts by asymptotes. That means that the lines x=4 and y=5 are lies that can not be crossed by this equation.

OpenStudy (anonymous):

Ok, so how would I solve this?

OpenStudy (e.mccormick):

Well, I would look at the standard graph of \(y=\dfrac{6}{h}\) and the existing asymptotes, then see what would be needed to shift the center to the new asymptotes.

OpenStudy (e.mccormick):

If you understand how and why it works, you can do it mathematically. Until you do, it is easier to look at it graphically: https://www.desmos.com/calculator/cun2ko8ds5 So play with that \(y=\dfrac{6}{x-h}+k\) form I mentioned until you get the graph to center over where the x=4 and y=5 lines cross.

OpenStudy (anonymous):

Ok, so

OpenStudy (anonymous):

y= 6/x-4 +5?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@e.mccormick ? Are you there?

OpenStudy (e.mccormick):

Odd, it was not updating my display and I had to refresh. I am going to assume you mean (x-4) as in all under the /, then yes.

OpenStudy (anonymous):

yep, thanks :)